# Scattering Gravitons off General Spinning Compact Objects to $\mathcal{O}(G^{2}S^{4})$

### v1

@@ Abstract
Introduction—Following the monumental detection of gravitational waves by the LIGO–Virgo–KAGRA collaboration Abbott et al. (2016, 2017), an unprecedented observational window into the universe has opened, enabling high-precision studies of binary systems composed of massive, compact astrophysical objects such as neutron stars and black holes. The advent of next-generation detectors demands corresponding theoretical progress—not only in precision, but also in our understanding of the fundamental aspects of classical two-body dynamics Punturo et al. (2010); Amaro-Seoane et al. (2017); Reitze et al. (2019); Borhanian and Sathyaprakash (2024); Pürrer and Haster (2020).
@@
Over the past several years, methods inspired by quantum field theory have driven remarkable advances in computing classical gravitational observables. Among these, the post-Minkowskian (PM) expansion—a perturbative expansion in Newton’s constant GG—has proven particularly powerful. For spinless binaries, substantial progress has been made using both on-shell scattering amplitude approaches Cheung et al. (2018); Kosower et al. (2019); Bern et al. (2019a, b); Cristofoli et al. (2019); Bjerrum-Bohr et al. (2020); Brandhuber et al. (2021); Bern et al. (2021a, 2022); Damgaard et al. (2023); Kosmopoulos and Solon (2024); Bern et al. (2025) and worldline methods Kälin and Porto (2020); Kälin et al. (2020, 2023); Dlapa et al. (2023, 2022); Mogull et al. (2021); Jakobsen et al. (2021, 2022a); Jakobsen (2023); Driesse et al. (2024a, b); He and Parra-Martinez (2025); Cheung et al. (2023, 2024). Including spin effects, however, is essential for accurately modeling astrophysical systems, motivating extensive efforts in both amplitude- and worldline-based formulations Bini and Damour (2017, 2018); Vines (2018); Vines et al. (2019); Guevara (2019); Guevara et al. (2019a); Chung et al. (2019); Arkani-Hamed et al. (2020); Guevara et al. (2019b); Chung et al. (2020a); Damgaard et al. (2019); Aoude et al. (2020); Chung et al. (2020b); Guevara et al. (2021); Bern et al. (2021b); Kosmopoulos and Luna (2021); Chen et al. (2022a, b); Febres Cordero et al. (2023); Bern et al. (2023, 2024); Menezes and Sergola (2022); Riva et al. (2022); Damgaard et al. (2022); Aoude et al. (2022a, b); Bautista et al. (2023a); Gonzo and Shi (2023); Aoude et al. (2023); Lindwasser (2024a); Brandhuber et al. (2024a); De Angelis et al. (2024); Aoude et al. (2024); Bohnenblust et al. (2024a); Gatica (2024); Cristofoli et al. (2024); Luna et al. (2024); Gatica (2023); Liu et al. (2021); Jakobsen et al. (2022b, c); Jakobsen and Mogull (2022, 2023); Jakobsen et al. (2023a, b); Heissenberg (2023); Lindwasser (2024b); Bautista et al. (2024); Cangemi et al. (2024); Brandhuber et al. (2024b); Chen and Wang (2025); Correia and Isabella (2024); Bhattacharyya et al. (2025); Alaverdian et al. (2025); Brandhuber et al. (2024c, d); Akpinar et al. (2025a, b); Bohnenblust et al. (2024b); Haddad et al. (2024); Bonocore et al. (2024); Akpinar et al. (2025c); Bohnenblust et al. (2025); Aoki et al. (2025a); Hoogeveen et al. (2025); Akpinar et al. (2025d); Aoki et al. (2025b); Aoude and Helset (2025).
@@
While a substantial body of existing work has concentrated on Kerr black holes at the first post-Minkowskian (1PM) order Guevara et al. (2019a); Chung et al. (2019); Bern et al. (2023); Aoude et al. (2022a, 2023); Bautista and Guevara (2019); Bautista et al. (2023a); Arkani-Hamed et al. (2021); Falkowski and Machado (2021); Chiodaroli et al. (2022); Cangemi et al. (2023), comparatively fewer studies have explored the dynamics of generic spinning compact objects Haddad (2023); Ben-Shahar (2024); Saketh and Vines (2022); Vazquez-Holm and Luna (2025). Moreover, the full classical gravitational Compton amplitude at the second post-Minkowskian (2PM) order has only very recently been computed for Schwarzschild black holes Bjerrum-Bohr et al. (2025); Brunello et al. (2025) [^1] The full classical spinless one-loop Compton for electrodynamics has been computed in Ref Brunello and De Angelis (2024)., and for Kerr black holes through the quadratic order in spin Akpinar et al. (2025c). In fact, generic spinning bodies at this order have also previously been studied, but only within the eikonal limit Chen et al. (2022c). In this paper, we report the full classical one-loop gravitational Compton amplitude describing the scattering of a graviton off a generic massive spinning compact object at 2PM, $\mathcal{O}(G^{2})$, including terms up to the quartic order in spin. This extends the recent 2PM results Bjerrum-Bohr et al. (2025); Akpinar et al. (2025c) to higher orders in spin, including spin-induced non-minimal couplings. From this result we extract the gauge-invariant scattering phase in momentum space—valid in the eikonal and wave regimes—whose position space representation in the eikonal regime serves as a generator of observables Alessio et al. (2025); Kim (2025); Haddad et al. (2025); Akhtar et al. (2025); Ivanov et al. (2025). Invoking the classical equivalence principle of general relativity, we extract the eikonal-regime graviton monopole contribution in position space and identify it with the dynamics of a massless scalar probe in a Kerr background for aligned-spin configurations. The monopole term—proportional to the squared Lorentz product of the graviton polarizations—encodes the spin-independent information of the massless field.
@@
The organization of the paper is as follows. As our construction requires the regularization of loops, we begin with an outline of the different dimensional regularization schemes considered throughout. Then, we overview the relevant tree-level amplitudes required for the generalized unitarity framework, focusing first on the Kerr dynamics. Following this, we describe the detailed computation of the classical one-loop Compton amplitude, providing a workflow agnostic as to whether the dynamics are minimal or non-minimal. In fact, several properties arising purely from quantum field theory considerations, such as Weinberg’s infrared factorization, serve as strong consistency checks of our results. Equipped with this knowledge, we obtain the classical one-loop Compton amplitude that describes the scattering of a graviton off a Kerr black hole through the quartic order in spin, thereby extending the results of Refs. Bjerrum-Bohr et al. (2025); Brunello et al. (2025); Akpinar et al. (2025c). In doing so, we find that the amplitude is dependent on the dimensional regularization schemes that we consider in this paper at the finite order of the dimensional regularization parameter, i.e. at $\mathcal{O}(\epsilon^{0})$, starting at the quadratic order in spin. Such contributions, however, do not play a role in the scattering phase, which follows by subtracting the tree iteration from the one-loop amplitude. Within the eikonal regime, we find that the scattering phase for the graviton monopole contribution exhibits the anticipated spin-shift symmetry through the quartic order in spin Chen et al. (2022c); Akpinar et al. (2025c). Furthermore, in position space we find an exact agreement with the dynamics of a massless scalar probe in a Kerr background for aligned-spin configurations. Incorporating spin-induced non-minimal couplings Haddad et al. (2024), we further calculate the non-minimal classical one-loop Compton amplitude through the quartic order in spin. As before, we obtain the momentum space scattering phase, valid in both the eikonal and wave regimes, as well as the eikonal-regime monopole contribution to the scattering phase in position space. To conclude, we summarize our work and discuss future directions that build upon these results.
@@
THV and DR produce identical results for the classical one-loop Compton, while CDR differs starting at the $\epsilon^{0}$ order for quadratic and higher orders in spin. This difference stems from explicit factors of $D$ appearing in contractions of the spin tensor in tree-level amplitudes. We will revisit this point later.
@@
Classical Compton—All relevant contributions to the classical one-loop Compton amplitude are fixed by the t-channel and s-channel cut configurations; see Fig. 1. Within the generalized unitarity framework, the product of tree-level amplitudes requires a summation over physical states that can cross the unitarity cut. For massive legs, no such summation arises since the heavy-mass expansion enforces that all massive matter propagators are on-shell. For the massless legs, however, one requires the $D$-dimensional transverse graviton projector
@@
Then, choosing the basis $\{v^{\mu},k^{\mu},q^{\mu},n^{\mu}=\epsilon^{\mu\nu\rho\sigma}q_{\nu}v_{\rho}k_{\sigma}\}$, we proceed by decomposing all remaining vectors onto this basis; see the supplemental material for more details. The coefficients of (10) for the classical minimally coupled one-loop Compton through the quartic order in spin are provided in the ancillary file ComptonAmpsAndObs.m. These coefficients indeed satisfy the infrared factorization, the expected consistency relations, as well as gauge invariance through the quartic order in spin.
@@
In light of this observation, an important question becomes: are physical observables affected by regularization scheme dependences? It has recently been pointed out in Ref. Bini et al. (2024) that the spinless waveform at one-loop is dimensional regularization scheme dependent. As our discussion has revealed that the cut box and cut bubble master coefficients are dimensional regularization scheme dependent in the presence of spin, paired with the fact that the one-loop Compton amplitude plays a pivotal role in loop-level computations of the waveform, this highlights one source of scheme dependence that must be tracked consitently. It would be interesting to see exactly how this dimensional regularization scheme dependence affects the one-loop waveform, but we leave such a tantalizing endeavor to future work. If we restrict ourselves to non-radiative observables, however, then any scheme dependence is expected to drop out. This is the focus of the next section.
@@
In this limit, the finite part coming from the bubble $d_{\mathrm{R}}$ is quantum suppressed as compared to the triangle $d_{\bm{q}}$, such that only the latter contributes in the classical limit. Indeed the coefficients in (22) exhibit spin-shift symmetry through the quartic order in spin [^3] Note that spin-shift symmetry is not present beyond the eikonal limit, nor is it present beyond the monopole contribution of the graviton.; as previously observed in Refs. Chen et al. (2022c); Akpinar et al. (2025c). Explicitly, the coefficients read:
@@
The relevant integrals to $n=4$ are in the supplemental material, while the monopole contribution to the scattering phase in position space can be found in the ancillary file ComptonAmpsAndObs.m. Note that we do not include the full momentum space scattering phase (which is also valid in the wave regime), as the coefficients of the classical one-loop Compton are readily found in the file.
@@
Beyond Kerr dynamics—To deform the classical one-loop Compton amplitude beyond the Kerr black hole case and thereby describe the dynamics of a generic spinning compact object, we must incorporate spin-induced non-minimal effects from the quadratic order in spin onwards. Our starting point is the bosonic worldline model of Ref. Haddad et al. (2024), which includes such deformations through linear-curvature and leading higher-curvature couplings. Within this framework, the worldline action describing a massive spinning point particle with spin-induced non-minimal couplings reads
@@
such that the dynamics preserves the spin-supplementary condition $v_{\mu}S^{\mu\nu}=0$ where $v^{\mu}=\dot{x}^{\mu}$. One can then go ahead and derive Feynman rules via a weak-field expansion $g_{\mu\nu}=\eta_{\mu\nu}+\kappa h_{\mu\nu}$ where $h_{\mu\nu}$ is graviton field. At the same time, one must perturb the trajectory and bosonic oscillators around their background values; see Ref. Haddad et al. (2024) for more details. Importantly, we should note that the Feynman rules obtained from the BUSY model is valid in four-dimensions only, meaning that for the non-minimal computation we employ the THV and DR schemes.
@@
Obtaining the non-minimal classical four-point Compton amplitude follows by construct all possible diagrams with two graviton emissions from the worldline. Specifically, we need the diagrams with no external deflections, and at most a single internal deflection. There are three contributions to this amplitude: the $t$-channel diagram, which comes from the contraction of the single graviton emission from the worldline with the three-graviton vertex; the diagrams involving propagating worldline and oscillator deflections; and contact terms coming from the double graviton emission from the worldline.
One must actually take care when using the tree level Compton acquired here within the generalized unitarity framework, since Ref. Haddad et al. (2025) pointed out that the cubic in spin and quartic in spin contributions do not satisfy GWI. Here we find that only the cubic in spin terms do not satisfy this property, unless we impose four-dimensional identities; practically speaking, this is done by substituting a general frame valid in four-dimensions to make this property manifest. It seems that the cubic in spin structures do not even manifestly satisfy the expected Ward identity (WI) unless the four-dimensional identities are imposed. Of course, if we wish to use this amplitude in the generalized unitarity framework, GWI (or even WI) must be manifest. The standard way to remedy this would be to impose all four-dimensional identities by choosing a basis. In practice, however, this is complicated by the external graviton legs, which lead to many different Levi-Civita structures and momentum dependent denominators. By diagnosing the amplitude in detail, we find that contributions proportional to $C_{\mathrm{BS^{3}}}$ break this property. To circumvent this problem we simply add a Gram determinant—which vanishes in four-dimensions—to restore the expected property, without affecting the amplitude itself, at least in four-dimensions [^5] This was determined by simply taking the difference of the non-minimal tree-level amplitude with Kerr values and minimal tree-level amplitude.. One can think of this as simply adding a zero in four-dimensions. Once this has been done, the WI and GWI are indeed manifestly restored. For total transparency, both versions of the amplitude can be found in the ancillary file ComptonAmpsAndObs.m.
@@
Using these non-minimal tree-level amplitudes as the ingredients of our generalized unitarity pipeline yields the non-minimal classical one-loop Compton. Importantly, we find that the usual Kerr values ($C_{\mathrm{ES}^{n}}=C_{\mathrm{BS}^{n}}=1$) reproduce the minimally coupled amplitude. A crucial and immediate check is the agreement of the cut box coefficient obtained via the $t$- and $s$-channel cut configurations. In fact, this is a non-trivial check since the non-minimal information is encoded very differently in the three- and four-point tree-level amplitudes. Moreover, we find that the amplitude satisfies the expected infrared factoization (12), relation between coefficients, and gauge invariance. On the note of scheme dependence, we indeed find that the cut box and bubble master values are scheme dependent, but this dependence drops out after substituting masters, since we only use the THV and DR schemes. Finally, we compute the momentum space scattering phase and extract the monopole contribution to the graviton in the eikonal regime, which we Fourier transform into position space. As expected, we find that spin-shift symmetry is only satisfied by the eikonal-regime monopole contribution in momentum space for the Kerr values of the Wilson coefficients Aoude et al. (2021); Kosmopoulos and Luna (2021); Chen et al. (2022c); Akpinar et al. (2025d). Finally, we compute the aligned-spin scattering angle.
As before, the non-minimal classical one-loop Compton through the quartic order in spin, as well as the monopole contribution to the scattering phase in position space and scattering angle, can be found in ComptonAmpsAndObs.m.
@@
Our construction here incorporates both minimal and spin-induced non-minimal interactions, thereby capturing spin-induced finite-size effects associated with generic compact objects beyond the Kerr solution. The resulting amplitude satisfies several non-trivial consistency checks, including the Weinberg infrared factorization, expected relations among coefficients, as well as gauge invariance. Interestingly, we found that the one-loop Compton seems to be dimensional regularization scheme dependent in the presence of spin, which we anticipate to be a source of scheme dependence appearing in waveform computations Bini et al. (2024). The one-loop Compton is indeed an essential ingredient in, for example, computations of the loop-level waveform Herderschee et al. (2023); Georgoudis et al. (2023); Brandhuber et al. (2023); Elkhidir et al. (2024); Bohnenblust et al. (2024a, 2025); Brunello and De Angelis (2024); Brunello et al. (2025).
To study non-radiative observables we extracted the scattering phase, which we found to be independent of the infrared divergence, the superclassical contributions, and the dimensional regularization scheme. In fact, this constitutes the momentum space scattering angle which is valid in the eikonal and wave regimes. By analyzing the monopole contribution of the graviton in the eikonal regime in position space—with and without non-minimal couplings—we found exact agreement with the dynamics of a massless scalar probe in a Kerr background, for Kerr values of Wilson coefficients. As anticipated, we found that this scattering phase exhibits spin-shift symmetry for Kerr values of the Wilson coefficients Chen et al. (2022c); Akpinar et al. (2025c). Finally, from this position space scattering phase we further extracted the aligned-spin scattering angle.
@@
Acknowledgements—I thank Rafael Aoude, Fabian Bautista, Graham R. Brown, Giacomo Brunello, Asaad Elkhidir, Riccardo Gonzo and Mao Zeng for many insightful discussions and comments on the draft. I am especially grateful to Kays Haddad for sharing private files for the Feynman rules of Ref. Haddad et al. (2024) and for many insightful discussions. I have extensively used the Mathematica package FeynCalc Shtabovenko et al. (2025) for some of these calculations. D.A. is supported by an STFC studentship.

### v2

@@ Abstract
Introduction—Following the monumental detection of gravitational waves by the LIGO–Virgo–KAGRA collaboration Abbott et al. (2016, 2017), an unprecedented observational window into the universe has opened, enabling high-precision studies of binary systems composed of massive, compact astrophysical objects like neutron stars and black holes. The advent of next-generation detectors demands corresponding theoretical progress—not only in precision, but also in our understanding of the fundamental aspects of classical two-body dynamics Punturo et al. (2010); Amaro-Seoane et al. (2017); Reitze et al. (2019); Borhanian and Sathyaprakash (2024); Pürrer and Haster (2020).
@@
Over the past several years, methods inspired by quantum field theory have driven remarkable advances in computing classical gravitational observables. Among these, the post-Minkowskian (PM) expansion—a perturbative expansion in Newton’s constant GG—has proven particularly powerful. For spinless binaries, substantial progress has been made using both on-shell scattering amplitude approaches Cheung et al. (2018); Kosower et al. (2019); Bern et al. (2019a, b); Cristofoli et al. (2019); Bjerrum-Bohr et al. (2020); Brandhuber et al. (2021); Bern et al. (2021a, 2022); Damgaard et al. (2023); Kosmopoulos and Solon (2024); Bern et al. (2025) and worldline methods Kälin and Porto (2020); Kälin et al. (2020); Dlapa et al. (2022a); Kälin et al. (2023); Dlapa et al. (2023a, b, 2022b); Mogull et al. (2021); Jakobsen et al. (2021, 2022a); Jakobsen (2023); Driesse et al. (2024a, b); He and Parra-Martinez (2025); Cheung et al. (2023, 2024). Including spin effects, however, is essential for accurately modeling astrophysical systems, motivating extensive efforts in both amplitude- and worldline-based formulations Porto (2006); Porto and Rothstein (2008); Porto (2016); Bini and Damour (2017, 2018); Vines (2018); Vines et al. (2019); Guevara (2019); Guevara et al. (2019a); Chung et al. (2019); Arkani-Hamed et al. (2020); Guevara et al. (2019b); Chung et al. (2020a); Damgaard et al. (2019); Aoude et al. (2020); Chung et al. (2020b); Guevara et al. (2021); Bern et al. (2021b); Kosmopoulos and Luna (2021); Chen et al. (2022a); Febres Cordero et al. (2023); Bern et al. (2023, 2024); Menezes and Sergola (2022); Riva et al. (2022); Damgaard et al. (2022); Aoude et al. (2022a, b); Bautista et al. (2023a); Gonzo and Shi (2023); Aoude et al. (2023); Lindwasser (2024a); Brandhuber et al. (2024a); De Angelis et al. (2024); Aoude et al. (2024); Bohnenblust et al. (2024a); Gatica (2024); Cristofoli et al. (2024); Luna et al. (2024); Gatica (2023); Liu et al. (2021); Jakobsen et al. (2022b, c); Jakobsen and Mogull (2022, 2023); Jakobsen et al. (2023a, b); Heissenberg (2023); Lindwasser (2024b); Bautista et al. (2024); Cangemi et al. (2024); Brandhuber et al. (2024b); Chen and Wang (2025); Correia and Isabella (2024); Bhattacharyya et al. (2025); Alaverdian et al. (2025); Brandhuber et al. (2024c, d); Akpinar et al. (2025a, b); Bohnenblust et al. (2024b); Haddad et al. (2024); Bonocore et al. (2024); Akpinar et al. (2025c); Bohnenblust et al. (2025); Aoki et al. (2025a); Hoogeveen et al. (2025); Akpinar et al. (2025d); Aoki et al. (2025b); Aoude and Helset (2025).
@@
While a substantial body of existing work has concentrated on Kerr black holes at the first post-Minkowskian (1PM) order Guevara et al. (2019a); Chung et al. (2019); Bern et al. (2023); Aoude et al. (2022a, 2023); Bautista and Guevara (2019); Bautista et al. (2023a); Arkani-Hamed et al. (2021); Falkowski and Machado (2021); Chiodaroli et al. (2022); Cangemi et al. (2023), comparatively fewer studies have explored the dynamics of generic spinning compact objects Haddad (2023); Ben-Shahar (2024); Saketh and Vines (2022); Vazquez-Holm and Luna (2025). Moreover, the full classical gravitational Compton amplitude at the second post-Minkowskian (2PM) order has only very recently been computed for Schwarzschild black holes Bjerrum-Bohr et al. (2025); Brunello et al. (2025) [^1] The full classical spinless one-loop Compton for electrodynamics has been computed in Ref Brunello and De Angelis (2024)., and for Kerr black holes through the quadratic order in spin Akpinar et al. (2025c). In fact, generic spinning bodies at this order have also previously been studied, but only within the eikonal limit Chen et al. (2022b). In this paper, we report the full classical one-loop gravitational Compton amplitude describing the scattering of a graviton off a generic massive spinning compact object at 2PM, $\mathcal{O}(G^{2})$, including terms up to the quartic order in spin. This extends the recent 2PM results Bjerrum-Bohr et al. (2025); Akpinar et al. (2025c) to higher orders in spin, including spin-induced non-minimal couplings. From this result we extract the gauge-invariant scattering phase in momentum space—valid in the eikonal and wave regimes—whose position space representation in the eikonal regime serves as a generator of observables Alessio et al. (2025); Kim (2025); Haddad et al. (2025); Akhtar et al. (2025); Ivanov et al. (2025). Invoking the classical equivalence principle of general relativity, we extract the eikonal-regime graviton monopole contribution in position space and identify it with the dynamics of a massless scalar probe in a Kerr background for aligned-spin configurations. The monopole term—proportional to the squared Lorentz product of the graviton polarizations—encodes the spin-independent information of the massless field.
@@
The organization of the paper is as follows. As our construction requires the regularization of loops, we begin with an outline of the different dimensional regularization schemes considered throughout. Then, we overview the relevant tree-level amplitudes required for the generalized unitarity framework, focusing first on the Kerr dynamics. Following this, we describe the detailed computation of the classical one-loop Compton amplitude, providing a workflow agnostic as to whether the dynamics are minimal or non-minimal. In fact, several properties arising purely from quantum field theory considerations, such as Weinberg’s infrared factorization, serve as strong consistency checks of our results. Equipped with this knowledge, we obtain the classical one-loop Compton amplitude that describes the scattering of a graviton off a Kerr black hole through the quartic order in spin, thereby extending the results of Refs. Bjerrum-Bohr et al. (2025); Brunello et al. (2025); Akpinar et al. (2025c). In doing so, we find that the amplitude is dependent on the dimensional regularization schemes that we consider in this paper at the finite order of the dimensional regularization parameter, i.e. at $\mathcal{O}(\epsilon^{0})$, starting at the quadratic order in spin. Such contributions, however, do not play a role in the scattering phase, which follows by subtracting the tree iteration from the one-loop amplitude. Within the eikonal regime, we find that the scattering phase for the graviton monopole contribution exhibits the anticipated spin-shift symmetry through the quartic order in spin Chen et al. (2022b); Akpinar et al. (2025c). Furthermore, in position space we find an exact agreement with the dynamics of a massless scalar probe in a Kerr background for aligned-spin configurations. Incorporating spin-induced non-minimal couplings Haddad et al. (2024), we further calculate the non-minimal classical one-loop Compton amplitude through the quartic order in spin. As before, we obtain the momentum space scattering phase, valid in both the eikonal and wave regimes, as well as the eikonal-regime monopole contribution to the scattering phase in position space. To conclude, we summarize our work and discuss future directions that build upon these results.
@@
The THV and DR schemes produce identical results for the classical one-loop Compton amplitude, while the CDR scheme differs starting at the $\epsilon^{0}$ order for quadratic and higher orders in spin. This difference stems from explicit factors of $D$ appearing in contractions of the spin tensor in tree- level amplitudes. We will revisit this point later.
@@
The Classical Compton Amplitude—All relevant contributions to the classical one-loop Compton amplitude are fixed by the t-channel and s-channel cut configurations; see Fig. 1. Within the generalized unitarity framework, the product of tree-level amplitudes requires a summation over physical states that can cross the unitarity cut. For massive legs, no such summation arises since the heavy-mass expansion enforces that all massive matter propagators are on-shell. For the massless legs, however, one requires the $D$-dimensional transverse graviton projector
@@
Then, choosing the basis $\{v^{\mu},k^{\mu},q^{\mu},n^{\mu}=\epsilon^{\mu\nu\rho\sigma}q_{\nu}v_{\rho}k_{\sigma}\}$, we proceed by decomposing all remaining vectors onto this basis; see the supplemental material for more details. The coefficients of (10) for the classical minimally coupled one-loop Compton amplitude through the quartic order in spin are provided in the ancillary file ComptonAmpsAndObs.m. These coefficients indeed satisfy the infrared factorization, the expected consistency relations, as well as gauge invariance through the quartic order in spin.
@@
In light of this observation, an important question becomes: are physical observables affected by regularization scheme dependences? It has recently been pointed out in Ref. Bini et al. (2024) that the spinless waveform at one-loop is dimensional regularization scheme dependent. As our discussion has revealed that the cut box and cut bubble master coefficients are dimensional regularization scheme dependent in the presence of spin, paired with the fact that the one-loop Compton amplitude plays a pivotal role in loop-level computations of the waveform, this highlights one source of scheme dependence that must be tracked consistently. It would be interesting to see exactly how this dimensional regularization scheme dependence affects the one-loop waveform, but we leave such a tantalizing endeavor to future work. If we restrict ourselves to non-radiative observables, however, then any scheme dependence is expected to drop out. This is the focus of the next section.
@@
In this limit, the finite part coming from the bubble $d_{\mathrm{R}}$ is quantum suppressed as compared to the triangle $d_{\bm{q}}$, such that only the latter contributes in the classical limit. Indeed the coefficients in (22) exhibit spin-shift symmetry through the quartic order in spin [^3] Note that spin-shift symmetry is not present beyond the eikonal limit, nor is it present beyond the monopole contribution of the graviton.; as previously observed in Refs. Chen et al. (2022b); Akpinar et al. (2025c). Explicitly, the coefficients read:
@@
The relevant integrals to $n=4$ are in the supplemental material, while the monopole contribution to the scattering phase in position space can be found in the ancillary file ComptonAmpsAndObs.m. Note that we do not include the full momentum space scattering phase (which is also valid in the wave regime), as the coefficients of the classical one-loop Compton amplitude are readily found in the file.
@@
Beyond Kerr dynamics—To deform the classical one-loop Compton amplitude beyond the Kerr black hole case and thereby describe the dynamics of a generic spinning compact object, we must incorporate spin-induced non-minimal effects from the quadratic order in spin onward. Our starting point is the bosonic worldline model of Ref. Haddad et al. (2024), which includes such deformations through linear-curvature and leading higher-curvature couplings. Within this framework, the worldline action describing a massive spinning point particle with spin-induced non-minimal couplings reads
@@
such that the dynamics preserves the spin-supplementary condition $v_{\mu}S^{\mu\nu}=0$ where $v^{\mu}=\dot{x}^{\mu}$. One can then go ahead and derive Feynman rules via a weak-field expansion $g_{\mu\nu}=\eta_{\mu\nu}+\kappa h_{\mu\nu}$ where $h_{\mu\nu}$ is graviton field. At the same time, one must perturb the trajectory and bosonic oscillators around their background values; see Ref. Haddad et al. (2024) for more details. Importantly, we should note that the Feynman rules obtained from this model is valid in four-dimensions only, meaning that for the non-minimal computation we employ the THV and DR schemes.
@@
Obtaining the non-minimal classical four-point Compton amplitude follows by constructing all possible diagrams with two graviton emissions from the worldline. Specifically, we need the diagrams with no external deflections, and, at most, a single internal deflection. There are three contributions to this amplitude: the $t$-channel diagram, which comes from the contraction of the single graviton emission from the worldline with the three-graviton vertex; the diagrams involving propagating worldline and oscillator deflections; and contact terms coming from the double graviton emission from the worldline.
One must actually take care when using the tree-level Compton amplitude acquired here within the generalized unitarity framework, since Ref. Haddad et al. (2025) pointed out that the cubic in spin and quartic in spin contributions do not satisfy GWI. Here we find that only the cubic in spin terms do not satisfy this property, unless we impose four-dimensional identities; practically speaking, this is done by substituting a general frame valid in four-dimensions to make this property manifest. It seems that the cubic in spin structures also do not satisfy the expected Ward identity (WI) unless the four-dimensional identities are imposed. Of course, if we wish to use this amplitude in the generalized unitarity framework, GWI (or even WI) must be manifest. The standard way to remedy this would be to impose all four-dimensional identities by choosing a basis. In practice, however, this is complicated by the external graviton legs, which lead to many different Levi-Civita structures and momentum dependent denominators. By diagnosing the amplitude in detail, we find that contributions proportional to $C_{\mathrm{BS^{3}}}$ break this property. To circumvent this problem we simply add a Gram determinant—which vanishes in four-dimensions—to restore the expected property, without affecting the amplitude itself, at least in four-dimensions [^5] This was determined by simply taking the difference of the non-minimal tree-level amplitude with Kerr values and minimal tree-level amplitude.. One can think of this as simply adding a zero in four-dimensions. Once this has been done, the WI and GWI are indeed manifestly restored. For total transparency, both versions of the amplitude can be found in the ancillary file ComptonAmpsAndObs.m.
@@
Using these non-minimal tree-level amplitudes as the ingredients of our generalized unitarity pipeline yields the non-minimal classical one-loop Compton amplitude. Importantly, we find that the usual Kerr values ($C_{\mathrm{ES}^{n}}=C_{\mathrm{BS}^{n}}=1$) reproduce the minimally coupled amplitude. A crucial and immediate check is the agreement of the cut box coefficient obtained via the $t$- and $s$-channel cut configurations. In fact, this is a non-trivial check since the non-minimal information is encoded very differently in the three- and four-point tree-level amplitudes. Moreover, we find that the amplitude satisfies the expected infrared factorization (12), relation between coefficients, and gauge invariance. On the note of scheme dependence, we indeed find that the cut box and bubble master values are scheme dependent, but this dependence drops out after substituting masters, since we only use the THV and DR schemes. Finally, we compute the momentum space scattering phase and extract the monopole contribution to the graviton in the eikonal regime, which we Fourier transform into position space. As expected, we find that spin-shift symmetry is only satisfied by the eikonal-regime monopole contribution in momentum space for the Kerr values of the Wilson coefficients Aoude et al. (2021); Kosmopoulos and Luna (2021); Chen et al. (2022b); Akpinar et al. (2025d). Finally, we compute the aligned-spin scattering angle.
As before, the non-minimal classical one-loop Compton amplitude through the quartic order in spin, as well as the monopole contribution to the scattering phase in position space and scattering angle, can be found in ComptonAmpsAndObs.m.
@@
Our construction here incorporates both minimal and spin-induced non-minimal interactions, thereby capturing spin-induced finite-size effects associated with generic compact objects beyond the Kerr solution. The resulting amplitude satisfies several non-trivial consistency checks, including the Weinberg infrared factorization, expected relations among coefficients, as well as gauge invariance. Interestingly, we found that the one-loop Compton amplitude seems to be dimensional regularization scheme dependent in the presence of spin, which we anticipate to be a source of scheme dependence appearing in waveform computations Bini et al. (2024). The one-loop Compton amplitude is indeed an essential ingredient in, for example, computations of the loop-level waveform Herderschee et al. (2023); Georgoudis et al. (2023); Brandhuber et al. (2023); Elkhidir et al. (2024); Bohnenblust et al. (2024a, 2025); Brunello and De Angelis (2024); Brunello et al. (2025).
To study non-radiative observables we extracted the scattering phase, which we found to be independent of the infrared divergence, the superclassical contributions, and the dimensional regularization scheme. In fact, this constitutes the momentum space scattering phase which is valid in the eikonal and wave regimes. By analyzing the monopole contribution of the graviton in the eikonal regime in position space—with and without non-minimal couplings—we found exact agreement with the dynamics of a massless scalar probe in a Kerr background, for Kerr values of Wilson coefficients. As anticipated, we found that this contribution to the scattering phase in momentum space exhibits spin-shift symmetry for Kerr values of the Wilson coefficients Chen et al. (2022b); Akpinar et al. (2025c). Finally, from the position space scattering phase we further extracted the aligned-spin scattering angle.
@@
Acknowledgements—I thank Rafael Aoude, Fabian Bautista, Graham R. Brown, Giacomo Brunello, Asaad Elkhidir, Riccardo Gonzo, Matteo Sergola and Mao Zeng for many insightful discussions and comments on the draft. I am especially grateful to Kays Haddad for sharing private files for the Feynman rules of Ref. Haddad et al. (2024) and for many insightful discussions. I have extensively used the Mathematica package FeynCalc Shtabovenko et al. (2025) for some of these calculations. D.A. is supported by an STFC studentship.